OpenStudy (anonymous):

lg (20x^2 + 100x + 120) = 0.05 x = ???

OpenStudy (kira_yamato):

\[20x^2 + 100x + 120 = 10^{0.05}\] \[20x^2 + 100x + 120 - 10^{0.05} = 0\] Use the quadratic formula to solve the equation.

OpenStudy (anonymous):

ohhh. pellet. My brain is retarded. I'm weird.

OpenStudy (anonymous):

\[20 x^2+100x+120\text{ =}\frac{5}{100} \]Is the preceding the equation?